OpenStudy (anonymous):

435,000 is what percent of 725,000.

OpenStudy (anonymous):

I dont know how to figure that out

OpenStudy (anonymous):

[(435,000)/(725,000)]*100% = 60%

OpenStudy (mathstudent55):

You can do it with a propoertion. 725,000 is to 100% as 435,000 is x. \(\dfrac{725,000}{100} = \dfrac{435,000}{x} \) Cross multiply and solve for x.